A string of length 10 𝑓𝑑. is raised at the middle to a distance of 1 𝑓𝑑, and then released. Describe the motion of the string assuming 𝑐 = 1𝑓𝑑. 𝑠 βˆ’1.Β 


1
Expert's answer
2021-11-24T05:04:31-0500


Equation of AO

"y-0= \frac{x}{5}\\
y= \frac{x}{5}"


Equation of AB

"y= \frac{x-10}{5-10}\\
y= \frac{10-x}{5}\\"


Hence the motion distributed by

"y_{tt}=y_{xx}\\
y_{x,t}=X_{x}*xT''\\
\frac{T''}{T}=\frac{X''}{X}= - λ^2\\
BVP\\
X'' + λ^2X=0\\
X(0)=X(10)=0; T''+λ^2T=0; T'(0)=0\\
X(x)= c_1 \cos λ x + c_2 \sin λ x; T(t)= c_3 \cos λ t + c_4 \sin λ t\\
For \space X(0)=0 \implies c_1=0, T'(0)=0 \implies c_4 =0\\
X (10)=0 \implies 10 λ= n \pi \implies λ = \frac{n \pi}{10} ; T_n= \cos (\frac{n \pi}{10}t)\\
X_n= \sin(\frac{n \pi}{10}x)\\
Y(x,t)=\sum_{n=1}^{\infin} B_n\sin(\frac{n \pi}{10}x)* \cos (\frac{n \pi}{10}t)\\
Y(x,0)= f(x)=\sum_{n=1}^{\infin} B_n\sin(\frac{n \pi}{10}x)\\
B_n=\frac{2}{10} \int_0^{10} f(x) \sin(\frac{n \pi}{10}x) dx\\
= \frac{1}{5}[ \int_0^{5} \frac{1}{5} x \sin(\frac{n \pi}{10}x) dx+ \int_5^{10} \frac{1}{5}(10- x) \sin(\frac{n \pi}{10}x) dx]\\
= \frac{8}{n^2 \pi^2}\sin(\frac{n \pi}{2})\\
\therefore Y(x,t)=\sum_{n=1}^{\infin} \frac{8}{n^2 \pi^2}\sin(\frac{n \pi}{2}) \sin(\frac{n \pi}{10}x)* \cos (\frac{n \pi}{10}t)\\"